Problem #1: A press brake is used to bend metal. Currently, the press brake your company owns has a market value of $50,000. The machine's operating cost is $20,000 per year. If the machine is kept for one year it will have a salvage value of $40,000. If the machine is kept for two years, it will have a salvage value $10,000. Your company could hire another company to bend metal for your parts. The other company quoted you at $42,000 per year. Assume a %10 interest rate. Should your company continue using the press brake or hire the other company. If you decide to hire the other company, when should you?
Answer #1

EUAC press brake for one year = 50000 * (A/P, 10%,1) + 20000 - 40000 * (A/F, 10%,1)

= 50000 * 1.1 + 20000 - 40000 * 1

= 35000

EUAC press brake for two year = 50000 * (A/P, 10%,2) + 20000 - 10000 * (A/F, 10%,2)

= 50000 * 0.576190 + 20000 - 10000 * 0.476190

= 44047.62

As EUAC of hiring outside company = 42000

EUAC of existing company is less for one year than the outside company, so company should keep press for one year